# Splat-Nav: Safe Real-Time Robot Navigation in Gaussian Splatting Maps

### v1

@@ III-C Collision Detection
Consider two ellipsoid $\mathcal{E}_{a},\mathcal{E}_{b}$ (with means $\mu_{a},\mu_{b}$ and covariances $\Sigma_{a},\Sigma_{b}$) and the concave scalar function $K:(0,1)\to\mathbb{R}$},
$K(s)=(\mu_{b}-\mu_{a})^{T}\Bigl[\frac{1}{1-s}\Sigma_{a}+\frac{1}{s}\Sigma_{b}\Bigr]^{-1}(\mu_{b}-\mu_{a}).$
Then, $\mathcal{E}_{a}\cap\mathcal{E}_{b}=\emptyset$ if and only if there exists $s\in(0,1)$ such that $K(s)>1$.
The proof for this is stated in Proposition 2 in Gilitschenski and Hanebeck [11]. Note that while $K$ is concave in $s$ (i.e., K-K is convex in $s$), it is not concave with respect to the means and variances.
Theorem 1 is a complete test that will always indicate whether two ellipsoids are in collision or not, however it is not computationally easy to solve. Although the existence of an $s$ can be solved by maximizing the concave function $K$ and asserting negativeness, such a routine can be expensive for the many checks that must happen between the robot body and the environment since it involves a matrix inversion. To remedy this issue, we eliminate the matrix inversion by operating in a shared basis for both $\Sigma_{A}$ and $\Sigma_{B}$.

### v2

@@ IV-B Collision Detection
Given two ellipsoids $\mathcal{E}_{a},\mathcal{E}_{b}$ (with means $\mu_{a},\mu_{b}$ and covariances $\Sigma_{a},\Sigma_{b}$) and the concave function $K:(0,1)\to\mathbb{R}$},
$K(s)=(\mu_{b}-\mu_{a})^{T}\Bigl[\frac{1}{1-s}\Sigma_{a}+\frac{1}{s}\Sigma_{b}\Bigr]^{-1}(\mu_{b}-\mu_{a}),$
$\mathcal{E}_{a}\cap\mathcal{E}_{b}=\emptyset$ if and only if there exists $s\in(0,1)$ such that $K(s)>1$.
Note that while $K$ is concave in $s$ (i.e., K-K is convex in $s$), it is convex with respect to the means and variances. Theorem 1 is a complete test that will always indicate whether two ellipsoids are in collision or not. We note, however, that solving the feasibility problem in Theorem 1 can be challenging, particularly in large-scale problems, where the feasibility problem has to solved for many pairs of ellipsoids with an associated matrix inversion procedure in each problem. In general, Gaussian Splatting environments consists of hundreds of thousands to tens of millions of Gaussians [8]. Consequently, we eliminate the matrix inversion by operating in a shared basis for both $\Sigma_{A}$ and $\Sigma_{B}$, for faster collision-checking, detailed in the following Proposition.